OpenStudy (anonymous):

Using the Pythagorean Theorem, find the length of a leg of a right triangle if the other leg is 8 feet long and the hypotenuse is 10 feet long. A. 6 B. 36 C. 12.81 D. 4√ 41

OpenStudy (anonymous):

10^2-8^2=x^2

OpenStudy (anonymous):

sqrt(100-64)=x sqrt(36)=x x=6

OpenStudy (anonymous):

A i agree

OpenStudy (anonymous):

Pythagoras theorem is c^2=a^2+b^2